Suppose you have celebrated your 19th birthday. A rich uncle set up a trust fund


for you that will pay you R100 000 when you turn 25. If the relevant discount rate is


11 percent, how much is the fund worth today?

1
Expert's answer
2022-03-15T16:07:38-0400

Pv=Fv(1+r)-n

where Pv is the present value

Fv is the future value

R is the rate of interest

N is the time

Pv=100000(1+0.11)-6

=100000*0.5346408

=53464.08

So the fund is worth R53464.08 today